OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1) .

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

the slope would be 4/3 ...i just dont know how to find the y intercept from this

terenzreignz (terenzreignz):

Still here? I'm glad you can find the slope, it's correct :) Now you need to find the equation of the line, but that's easy. When you have a slope \(\color{green} m\) and a point \((\color{red} a \ , \ \color{blue}b)\) Then the equation of the line is given by... \[\Large y = \color{green}m(x-\color{red}a)+ \color{blue} b\] You actually have two points here, you may use any one of them for your point \((\color{red} a \ , \ \color{blue}b)\)

OpenStudy (anonymous):

y=4/3(x+1)-3 y=4/3x + 4/3-3 y=4/3x-5/3 is this correct?

terenzreignz (terenzreignz):

\[\Large y = \frac43x - \frac53\] Yes, brilliant :) Now to get the y-intercept, simply set x = 0 and solve for y. Well done, there ;)

OpenStudy (anonymous):

y intercept = -5/3 and thanks !!(:

terenzreignz (terenzreignz):

No problem :)